# Solving Unbalanced Optimal Transport on Point Cloud by Tangent Radial Basis Function Method

### v1

@@ Abstract
1 Introduction.
@@ 5.1 Numerical examples with exact solutions
First, we solve a classical Poisson equation using the TRBF method to test its accuracy. Let the computational domain be $(t,x)\in[0,1]\times[0,1]$. And considering problems with exact solutions to $u_{e}=\cos(\pi t)\cos(\pi x)$:
$u_{tt}+\triangle u=$ $-2\pi^{2}\cos(\pi t)\cos(\pi x),$
$u_{t}=$ $-\sin(\pi t)\cos(\pi x),$
$u_{x}=$ $-\cos(\pi t)\sin(\pi x).$
Table 5.1.1: Error of the TRBF method
@@
8 8 1.088e-1 - 1.256e-1 -
16 16 3.548e-2 1.61 4.441e-2 1.49
32 32 1.450e-2 1.29 1.879e-2 1.24
64 64 6.579e-3 1.14 8.665e-3 1.11
128 128 3.136e-3 1.06 4.162e-3 1.05
256 256 1.531e-3 1.03 2.040e-3 1.02
From Table 5.1.1, we can find that as the mesh is refined in time and space, the results given by the TRBF method converge to the true solution approximately on the order of 1.
Next, we solve a OT problem on a 1D region with Gaussian distribution. Since this OT problem has exact solution, we can use it to verify the accuracy of the proposed method. The Gaussian functions used are defined as follows:
@@
Then we choose $\rho_{0}(x)=\rho_{G}(x,0.45,0.05)$ as the initial state and $\rho_{1}(x)=\rho_{G}(x,0.55,0.05)$ as the termination state. We compare the exact and computed wasserstein-fish-rao metric in Table 5.1.2.
Table 5.1.2: ADMM TRBF method for 1D Gaussian.
@@
8 8 0.01 0.0240856
16 16 0.01 0.0118708
32 32 0.01 0.0103813
64 64 0.01 0.0101568
128 128 0.01 0.0100775
As shown in Table 5.1.2, $W_{WFR-C}$ converges to the exact Wasserstein-Fisher-Rao metric, $W_{WFR-E}$ approximately in first order with respect to $\Delta t$ and $h$.
@@ 5.2 Unbalanced optimal transport on smooth surfaces
Figure 5.2.1: SOT examples ($\beta=1$).
@@
Figure 5.2.2: SUOT examples ($\beta=1.5$).
@@
Figure 5.2.4: SUOT test on point cloud for $S1$ and $S2$.
@@ 5.3 Unbalanced optimal transport on general point cloud
Figure 5.3.1: SUOT test on general surfaces ($\beta=1.5$).
@@ Acknowledgement
This work is supported by the National Natural Science Foundation of China (grants 12301538 and 92370125) and the National Key R&D Program of China (grant 2021YFA1001300).

### v2

@@ Abstract
{highlights}
This work provides a novel method to solve unbanlanced optimal transport on point cloud. By combining the dynamic formulation and ADMM, the key ingredient of the computation of UOT problem on point cloud becomes to efficiently solve a Poisson equation.
The Poisson equation is solved by a meshless tangent radial basis function method. The method only requires points and normal vectors, which significantly simplify the computation.
1 Introduction
@@ 5.1 Numerical examples with exact solutions
First, we solve a classical Poisson equation using the TRBF method to test its accuracy. Let the computational domain be $(t,x)\in[0,1]\times[0,1]$. The equation is
$u_{tt}+\triangle u=$ $-8\pi^{2}\cos(2\pi t)\cos(2\pi x),$ $(t,\boldsymbol{x})\in(0,1)\times(0,1),$
$u_{t}=$ $0,$ $(t,\boldsymbol{x})\in\{0,1\}\times[0,1],$
$u_{x}=$ $0,$ $(t,\boldsymbol{x})\in[0,1]\times\{0,1\}.$
with the exact solution
$u_{e}=\cos(2\pi t)\cos(2\pi x).$
Table 5.1.1: The spatiotemporal convergence order of the TRBF method error.
@@
8 8 2.167e-2 - 2.846e-2 -
16 16 4.042e-3 2.20 5.577e-3 2.13
32 32 9.088e-4 2.05 1.253e-3 2.05
64 64 2.143e-4 2.03 2.981e-4 2.02
128 128 5.209e-5 2.01 7.276e-5 2.01
256 256 1.278e-5 2.01 1.790e-5 2.01
In Table 5.1.1, We can see that the convergence rate is second order as spatial and time step go to zero. The similar result was also reported in [30].
Regarding the convergence of ADMM method, we compute an example of optimal transport between 1D Gaussian distributions. The Gaussian distribution used are given as follows:
@@
Then we choose $\rho_{0}(x)=\rho_{G}(x,0.45,0.05)$ as the initial state and $\rho_{1}(x)=\rho_{G}(x,0.55,0.05)$ as the termination state.
Table 5.1.2: Errors of costs for 1D OT example.
@@
8 8 $1.41\times 10^{-2}$
16 16 $1.87\times 10^{-3}$
32 32 $3.81\times 10^{-4}$
64 64 $1.57\times 10^{-4}$
128 128 $7.75\times 10^{-5}$
As shown in Table 5.1.2, the error of cost also converges to zero, however the rate is reduced to first order.
@@ 5.2 Unbalanced optimal transport on smooth surfaces
Figure 5.2.1: SOT examples ($\beta=1$) and the calculation times are 442s, 156s, 227s, 278s and 186s respectively.
@@
Figure 5.2.2: SUOT examples ($\beta=1.5$) and the calculation times are 450s, 160s, 236s, 287s and 192s respectively.
@@
Figure 5.2.4: SUOT test on point cloud for $S1$, $S2$ and the calculation times are both 452s.
@@
In addition, we compare the results on the torus with finite element based method (GrSfem) [23] and a neural network based method (NN)[32]. The results are shown in Fig.5.2.6.
      NN
      GrSfem
      TRBF
(a) $\rho(0,\boldsymbol{x})$
(b) $\rho(0.25,\boldsymbol{x})$
(c) $\rho(0.5,\boldsymbol{x})$
(d) $\rho(0.75,\boldsymbol{x})$
(e) $\rho(1,\boldsymbol{x})$
Figure 5.2.6: Compare with different methods on torus. The first line: method based on neural network; The second line: surface finite element method based on gradient recovery; Last line: our tangent radial basis function method.
It can be found that the results obtained by our method are almost consistent with those obtained by GrSfem. The accuracy of the neural network method is relatively low.
Finally, we compare the computational time and memory usage with GrSfem [23] and NN [32] on a sphere OT example.
Table 5.2.1: Memory usage (MB) with different size of point cloud.
$N_{n}$ 390 1806 3934 5878 11406
NN(cpu) 51237 53331 98597 84969 144085
NN(gpu) 1169 5228 10878 16202 32379
GrSfem(cpu) 38 223 559 886 1905
TRBF(cpu) 191 216 371 546 1400
Table 5.2.2: CPU time (s) with different size of of point cloud.
$N_{n}$ 390 1806 3934 5878 11406
NN (GPU) 10422 13578 18931 31746 48493
GrSfem 14 27 48 65 119
TRBF 93 442 1078 2146 3934
The maximum memory usage and computational time are listed in Table 5.2.1 and Table 5.2.2 respectively. Neural network based methods use much larger memory than finite element method or the proposed TRBF method. The computational time is also much longer. The memory usage of FEM and TRBF method are comparable. While the computational time of FEM method is much less than TRBF method. Notice that FEM requires mesh information and the mesh generation is not included in Table 5.2.2. Taking advantage of the mesh structure in FEM, The coefficient matrix is much sparser than that in TRBF method such that the assembling of coefficient matrix is much faster and the linear system is also much easier to solve.
@@ 5.3 Unbalanced optimal transport on general point cloud
Figure 5.3.1: SUOT test on general surfaces ($\beta=1.5$) and the calculation times are 1249s and 454s respectively.
@@ Acknowledgement
This work is supported by the National Natural Science Foundation of China (grants 92370125, 12301538) and the National Key R&D Program of China (grant 2021YFA1001300).